OpenStudy (anonymous):

2y+3=-1/3(x-2) in standard form

OpenStudy (anonymous):

include steps

OpenStudy (anonymous):

Standard form is y=mx+b First simplify the equation: 2y+3= -1/3(x-2)

OpenStudy (anonymous):

So combine like terms: 2y+3 = -1/3x +2/3 2y=-1/3x-7/3 To get rid of fractions, multiply by 3: 6y = -x - 7

OpenStudy (anonymous):

then divide it all by 6: y=-1/6x-7/6

OpenStudy (anonymous):

you do realize standard form is ax+by=c that is slope-intercept form you did

OpenStudy (anonymous):

Oh shoot, yea i forgot.. I'm old lol. Hang on, gimme a chance

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2y+3= -1/3(x-2) ax+by=c 2y+3=-1/3x+2/3 2y+1/3x=-7/3 To get rid of fractions, multiply by 3: 6y+x=-7

OpenStudy (anonymous):

or... x+6y = -7

OpenStudy (anonymous):

thxxxxx

OpenStudy (anonymous):

You're welcome. Medals are appreciated ^^